calculate the angles which the angular momentum vector forms with the z-axis in some 3 states of a particle on a sphere described by spherical harmonics Y2,2, Y2,-1, Y2-0
Answer #1

Y2,2 = (15/32\pi)1/2 sin2\Theta e-2i\Phi

Y2,-1 = (15/8\pi)1/2 cos\Theta sin\Theta ei\Phi

Y2,0 = (5/16\pi)1/2 (3cos2\Theta - 1)